OpenStudy (anonymous):

-5/6 - 1/3 = ? A) -1/2 B) 1 1/6 C) 1/3

OpenStudy (anonymous):

b

OpenStudy (anonymous):

@pengembara_bumi1 you sure??

OpenStudy (anonymous):

@Mertsj is he correct?

OpenStudy (mertsj):

What is the common denominator?

OpenStudy (anonymous):

These are fractions. You must first make their denominators the same. The first fraction has a denominator of 6, the second one has a denominator of 3. We notice that 3 and 6 have a common multiple, which is 6. So we can multiply the denominator of 1/3 by 2. But remember, whatever you do to the denominator, you must also do to the numerator. So if we multiply the denominator of 1/3 by 2, we must multiply its numerator by 2 as well. So this is what it will look like:\[\bf -\frac{5}{6}-\frac{1}{3}=-\frac{5}{6}-\frac{1 \ \ (\times 2)}{3 \ \ (\times2)}=-\frac{5}{6}-\frac{2}{6}\]Now that both fractions have the same denominator, we can subtract the numerators. That's how addition/subtract with fractions works. You first make the denominator the same and whatever you do to the denominator you must also do to the numerator. After that you just add/substract the top numbers depending on what you have been asked to do.\[= \bf \frac{-5}{6}-\frac{2}{6}=\frac{-5-2}{6}=-\frac{ -7}{ 6 }=-1 \frac{1}{6}\]

OpenStudy (anonymous):

@genius12 so its -1 1/6 ok now I know the last problem I wrote they gave me the wrong answer lol

OpenStudy (anonymous):

Yes the key here is to find the common denominators. Once you do that, just subtract the tops of both fractions and that's it.

OpenStudy (anonymous):

u need to make sure the base (denominator) is same :)